OpenStudy (anonymous):

(x^2-2x+4/x^2-5x+6)divided by(x^3+8/x^2-9)

OpenStudy (anonymous):

\[[(x^2-2x+4)\div(x^2-5x+6)] \div[(x^3+8/x^2-9)]\]

OpenStudy (anonymous):

Multiply the second part by the numerator and simplify.

OpenStudy (anonymous):

I factored down to the following: \[[(x-2)(x-2)/(x-6)(x-1)]*[(x-3)(x+3)/(x+2)(x^2-2x+4)]\]

OpenStudy (anonymous):

But there are no like terms, so do you just combine all of the factors into one problem? I got this as my answer so far: \[(x-2)^2(x-3)(x+3)/(x-1)(x-6)(x+2)(x^2-2x+4)\]

OpenStudy (anonymous):

You flip the second part, then multiply them out. Factoring is unnecessary.

OpenStudy (anonymous):

I did flip the terms, but the bases aren't the same.

OpenStudy (anonymous):

You're multiplying. It doesn't matter if the bases are the same. That only matters with addition and subtraction of fractions.

OpenStudy (anonymous):

So my answer is still technically right, I just have to multiply everything out?

OpenStudy (anonymous):

Yes, but I wouldn't.

OpenStudy (anonymous):

Leave it unfactored, flip it, and multiply through.

OpenStudy (anonymous):

But why include fractions that contain the difference and sums of two squares and cubes if you weren't supposed to factor them?

OpenStudy (anonymous):

Because you're not solving for the roots of x. You'd have to multiply it all out all over again anyway, and make things VERY difficult for yourself. Itd be redundant.

OpenStudy (anonymous):

Multiplying all of that out would take an entire page, and isn't factoring simplifying a problem? My instructions say to simplify completely.

OpenStudy (anonymous):

(didn't include that above, which is my bad)

OpenStudy (anonymous):

You have to factor AFTER simplifying, and AFTER multiplying... Otherwise nothing is simplified, and it's redundant and unnecessary.

OpenStudy (anonymous):

That and you didn't factor correctly, so I'd want to avoid that.

OpenStudy (anonymous):

So you are saying I should take \[(x^2-2x+4)(x^2-9)\] and so on?

OpenStudy (anonymous):

Yes. Then post what you get.

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

The numerator should be x^4 -2x^3 + 4x^2 -9x^2 +18x -36 simplify and factor that

OpenStudy (anonymous):

\[x^4-2x^3-5x^2-18x-36\] for above then?

OpenStudy (anonymous):

I had -9x^2+4x^2=-5x^2. Does that check out with you?

OpenStudy (anonymous):

Numerator yes. I'm not sure what you mean by the second post though

OpenStudy (anonymous):

For the numerator, it is fine though. That gives me plenty to work with, thank you again :D

OpenStudy (anonymous):

\[\left(\frac{4-2 x+x^2}{(-3+x) (-2+x)}\right)\left(\frac{(-3+x) (3+x)}{(2+x) \left(4-2 x+x^2\right)}\right)=\frac{3+x}{-4+x^2} \]

OpenStudy (anonymous):

Left Side = Right Side when both are evaluated with x=19